# Conditional and Modal Reasoning in Large Language Models

### v2

@@ 4 Results
GPT-4 Turbo (2024-04-09) 98.5 +0.2 +0.8
GPT-4 Turbo (1106) 96.5 +1.0 +2.7
GPT-4 (0613) 95.6 +0.8 +3.1
Gemini 1.5 Pro 92.1 +3.8 +6.7
GPT-4 (0314) 91.7 -1.9 +6.9
GPT-4o (2024-05-13) 87.9 +1.3 +9.6
Claude 3 Opus 82.9 +0.8 +14.0
Gemini 1.5 Flash 82.1 +0.4 +8.8
Llama 3 Instruct 70B 77.1 +1.7 +13.1
Mixtral 8x7B 75.8 +2.7 +12.9
Claude 3 Sonnet 74.6 +0.8 +10.2
@@
Claude 3 Haiku 69.0 -3.1 +9.6
Phi-2 60.0 +3.1 +16.7
Mistral 7B 59.8 -1.7 +10.2
Code Llama 13B 59.0 -1.0 +10.8
Code Llama 7B 57.5 -0.2 +11.5
GPT-3.5 Turbo (0613) 56.3 -1.0 +10.4
Code Llama 34B 56.3 -2.7 +9.8
GPT-3.5 Turbo (0125) 56.0 -3.8 +4.6
Llama 2 Chat 13B 55.2 -3.1 +3.8
Llama 2 Chat 70B 55.2 -5.2 +2.5
Llama 2 Chat 7B 55.0 -5.0 +0.2
GPT-3.5 Turbo (1106) 54.0 -4.2 -2.1
Yi Chat 34B 45.6 +4.2 +5.8
@@ Acknowledgments
For helpful feedback, we thank Dominic Hughes, Hailey Schoelkopf, and Shane Steinert-Threlkeld.

### v3

@@ 4 Results
GPT-4 Turbo (2024-04-09) 98.5 +0.2 +0.6
GPT-4 Turbo (1106) 96.5 +0.0 +2.1
GPT-4 (0613) 95.6 +0.2 +3.1
Gemini 1.5 Pro 92.1 -8.3 +6.7
GPT-4 (0314) 91.7 +5.8 +7.7
GPT-4o (2024-05-13) 87.9 +1.3 +9.0
Claude 3 Opus 82.9 +6.9 +14.6
Gemini 1.5 Flash 82.1 -3.3 +8.8
Llama 3 Instruct 70B 77.1 +5.4 +13.1
Mixtral 8x7B 75.8 -0.4 +12.9
Claude 3 Sonnet 74.6 +4.0 +10.2
@@
Claude 3 Haiku 69.0 -3.1 +7.7
Phi-2 60.0 -7.7 -8.1
Mistral 7B 59.8 -4.6 +9.2
Code Llama 13B 59.0 -9.0 +7.7
Code Llama 7B 57.5 +5.6 +0.2
GPT-3.5 Turbo (0613) 56.3 +1.7 +13.8
Code Llama 34B 56.3 -6.5 +4.4
GPT-3.5 Turbo (0125) 56.0 +2.1 +22.5
Llama 2 Chat 13B 55.2 +2.1 +0.0
Llama 2 Chat 70B 55.2 -3.1 +3.8
Llama 2 Chat 7B 55.0 -5.2 -3.5
GPT-3.5 Turbo (1106) 54.0 -0.4 +15.8
Yi Chat 34B 45.6 +4.4 +20.4
@@ Acknowledgments
For helpful feedback, we thank Dominic Hughes, James Kirkpatrick, Hailey Schoelkopf, and Shane Steinert-Threlkeld.